OpenStudy (anonymous):

find the point of intersection of the 2 lines and describe how you did it... y=2x and y=-1/2x+2

OpenStudy (anonymous):

im pretty sure you use substitution or elimination method but im not sure

OpenStudy (anonymous):

You can just set the two equations equal: 2x=(-1/2)x+2 Solve for x (5/2)x=2 X=4/5 y=2x=2(4/5)=8/5

OpenStudy (abstracthaven):